OpenStudy (anonymous):

x^2+y^2+4y=0 graph find the center and radius

OpenStudy (anonymous):

x^2 + (y+2)^2 -4=0 x^2 + (y+2)^2=2^2

OpenStudy (anonymous):

so, the radios r=2 and the center: (0,-2)

OpenStudy (anonymous):

this one is tricky. the general formula for a circle is \[(x - h)^{2} + (y - k)^{2} = r ^{2}\] so, you have to make the equation like this one. (x - 0)^2 + (Y^2 + 4y) = 0 the problem is in Y. so you have to complete the square. (x-0)^2 + (y +2)^2 = 4 the center is (0,-2) and the radius is 2.